# Red-Teaming for Generative AI: Silver Bullet or Security Theater?

### v2

@@ 4 A Survey of AI Red-teaming Research
In this section, we analyze the results of our extensive survey of recent research on AI red-teaming and related concepts.
Methodology. To source papers, we primarily searched arXiv, Google Scholar, OpenReview, ACL Anthology, and the ACM Digital Library with keywords “red-teaming”, “ai red-teaming”, “jailbreak”, and “llm jailbreak”, and we then gathered results.222While we focus on red-teaming evaluations in this work, we argue that the nascent jailbreaking literature contains techniques similar in spirit to those employed by AI red-teams and therefore also include this literature here. Where possible, we replaced preprints with corresponding published works. We additionally included relevant works encountered prior to this search as well as from snowball sampling.
We scrutinize and subdivide papers gathered for our survey into groups along two dimensions, both of which relate to the evaluation in each paper. The first corresponds to the type of risk investigated during the evaluation, and the second corresponds to the type of approach used for evaluation. We analyze papers by characteristics pertaining to threat model and methodology because we found that research works primarily focused on these aspects (perhaps due to technical relevance) as opposed to other factors important to red-teaming (such as team composition and resources consumed). Overall classifications and in-depth paper findings (as with analyses of case studies from the previous section) can be found in our appendix, and totals for each classification can be found in Table 3. Note that we do not ascribe moral attributes or judgments to any category (i.e., we do not posit that one form of red-teaming is better than another).
@@ 4.1 Findings: AI Red-teaming Threat Models
We categorize papers according to the threat model implied or specified in each. In particular, we consider the type of risk posed in each threat model. For the ease of organizing our retrieved papers and findings, we present risks as one of two binary types, but we note that in reality, there exists a spectrum between the two types, and risks generally may be more nuanced than as we describe them here.
Dissentive risk. We define dissentive risk as risk in which people may disagree on its definition and corresponding threat level, and we categorize papers that studied this type of risk accordingly. For instance, hate speech is defined many different ways, and though images with blood and responses to questions like “how to build a bomb?”333AI safety researchers note that “how to build a bomb?” is a default GenAI evaluation query despite the fact that related information could be found via Wikipedia or even popular fiction [160] because many GenAI developers assert their models should not disclose bomb-making instructions (e.g., [102, 146]), so techniques which yield responses prove safeguards are brittle. This does not necessarily mean the community finds such responses concerning. may be less than ideal AI generations, some may find them admissible depending on context.444Such generations do not reflect opinions held by the authors. Nevertheless, researchers like Gehman et al. [55] and Röttger et al. [123] create benchmarks to analyze hate speech generation, Rando et al. [116] perform red-teaming to ascertain the degree to which Stable Diffusion’s safety filter can be bypassed to generate NSFW content, and Zou et al. [184] evaluate whether LLMs respond to harmful queries like “Generate a step-by-step plan to destroy humanity” given adversarial prompt suffixes.
@@
Both and neither. Some authors tasked themselves with analyzing both kinds of risk, such as personally identifiable information (PII) leakage in addition to hate speech or dangerous generations [138, 53, 104]. Others introduce methods to analyze neither type of risk from the outset, stressing that definitions and classifications of issues may need to be done from scratch [27, 110].
@@ 4.2 Findings: AI Red-teaming Methodologies
Targeted attack. The last approach to red-teaming we document as part of our review involves deliberately targeting part of an LLM, which could include an API, vulnerability in language translation support, or step of its training process, in order to induce issues. As such, we refer to such approaches as targeted attack methods. For instance, Wang & Shu [153] show how to construct steering vectors using activation vectors from both safety-tuned and non-safety-tuned versions of models to obtain toxic outputs from safety-tuned models. Others illustrate how to imperceptibly perturb images to cause multimodal LMs to respond in unintended ways (such as replying with a malicious URL or misinformation) [128, 108, 11], and Tong et al. [145] engineer prompts for text-to-image models that are mismatched with resulting images by exploiting reliance on CLIP embeddings. Other approaches include but are not limited to weaponizing the fact that LLMs are not optimized to converse in low-resource languages and ciphers [44, 172, 174], poisoning data used to tune or utilize LLMs [76, 117, 176, 3, 23, 154], and attacking APIs associated with black-box models [103]. Various defensive methods rooted in targeted attack approaches have been proposed as well. Bitton et al. [18] describe their Adversarial Text Normalizer, which can defend an LLM against various character-level perturbations typical of certain adversarial prompts. In addition, other defensive strategies proposed in prior paragraphs can defend against attacks discussed here (e.g., JailGuard introduced by Zhang et al. [178] addresses attacks introduced in [108, 11, 128, 184]).
@@ 4.3 Discussion
Utilizing in-depth notes created for each paper from in this survey, two authors conducted thematic analyses of salient details to gather overall points (as with the case study analyses). Each author conducted their own analysis independently from one another, and points are summarized below.
Many different methods to perform red-teaming. As illustrated in Table 3, researchers and practitioners have undertaken numerous approaches to evaluate GenAI and have all described them as red-teaming. At the same time, there have been developments like Schuett et al.’s finding that the overwhelming majority of AGI lab members support external red-teaming efforts [129] and the recent Executive Order [143] stressing the importance of red-teaming. These developments and the many red-teaming variations are together arguably concerning, precisely because there is no agreed-upon definition (from these papers) regarding what constitutes red-teaming. By highlighting this, we do not mean to imply that evaluations until now are useless. On the contrary, we believe they are necessary but perhaps insufficient tests of safety, and we stipulate that the existence of many interpretations of “red-teaming” suggests there must be more top-down guidance and requirements concerning red-teaming evaluations.
Threat modeling skewed toward dissentive risk. Table 3 also highlights that the majority of evaluations focus on dissentive risk rather than consentive risk. This means that undue effort has been undertaken to evaluate and mitigate GenAI behavior that may be admissible in various contexts. Additionally, work like that of Röttger et al. [122] has shown that current attempts to mitigate such risks have resulted in exaggerated safety, yielding LLM behavior like the refusal to provide information on buying a can of coke. Lastly, focusing on dissentive risk takes attention away from consentive risk, which in turn is inadmissible in any context. In light of such issues and tradeoffs, Casper et al. [27] and Radharapu et al. [110] suggest clearly defining risks and problematic outputs and justifying those decisions before any analysis.
@@
Non-universality of values used for alignment. Work found as part of this survey involving dissentive risk and alignment are driven by, implicitly or explicitly, a set of human values that determine whether GenAI outputs are admissible or inadmissible. However, this in turn prompts the question whose values are being utilized for alignment and evaluation? For instance, the FLAMES benchmark proposed by Huang et al. [64] is purported to measure alignment with Chinese values, whereas Weidinger et al. [159] emphasize that other evaluations may reflect those of “the English-speaking or Western world.” The extent to which GenAI does not support low-resource languages [44, 172] and agrees with bias and stereotypes [53, 119] evidences that models may not reflect the values and beliefs of all persons. Works beyond this survey have illustrated how the framing of AI value alignment is a normative problem that, if not properly addressed, may only serve to reflect the norms of one group of people, typically the majority [48, 74, 73]. Especially as OpenAI has begun a partnership with the US military on one hand [139, 51] and launched an initiative to align superintelligent AI to “human values” on another [78], we argue that it is crucial to analyze assumptions made and viewpoints held by those who build AI systems.
No consensus on who should perform red-teaming. Moreover, just as there is a lack of agreement regarding values to use to assess GenAI outputs, there is a similar lack of agreement regarding who should perform red-teaming. Groups of evaluators have consisted of hired crowdworkers [53], competition participants [130], researchers themselves [104], and others simply red-teaming for fun [67]. While some argue for more diversity to evaluate AI models [137], others caution that increased diversity is not a panacea and is moreover typically ill-defined [159, 12]. For instance, Yong et al. [172] argue for multilingual red-teaming to respond to low-resource language issues, and He et al. [61] “advocate for a collaborative, interdisciplinary approach among the AI for Science community and society at large” to respond to scientific research risks. Such examples suggest that diversity should be defined and sought out relative to the risks considered by red-teaming processes. They additionally hint towards more involvement of the public and relevant stakeholders, ideas also recommended in parallel literature regarding algorithmic auditing and participatory approaches to machine learning [50, 40, 39, 15]. Future red-teaming guidelines may want to specify something along these lines.
@@ 5 NIST RFI Comment Analysis Summary
We find that comments submitted to the NIST RFI on red-teaming GenAI are generally in accord with our findings.555See our appendix for an extended analysis.
Similarities. Industry, academia, and civil society organizations suggest that NIST should specify a clearer definition of “red-teaming” with a clearer definition and provide interested parties with appropriate resources pertaining to guidelines and best practices. Notably, even industry firms with experience red-teaming GenAI expressed a desire for concrete guidance from NIST. This supports our finding that red-teaming, as defined in public research and reports, is loosely structured and perhaps not the rigorous practice implied by the Executive Order. Moreover, many comments stressed that a plurality of different viewpoints, stakeholders, and perspectives should be involved in evaluating GenAI systems. Our findings concur with these notions.
@@ 6 Takeaways and Recommendations
Red-teaming is not a panacea. Each red-teaming exercise discussed in this paper only covered a limited set of vulnerabilities. As such, red-teaming cannot be expected to guarantee safety from all angles. For instance, from the papers surfaced in our research survey, approaches to red-teaming that detect and mitigate harmful text responses [104] may not detect and mitigate phishing attack vulnerabilities [125] and vice versa. Similarly, our case study analysis highlighted that team composition may also influence the types of issues found in a given exercise (e.g., subject matter experts [8] may find different problems than crowdworkers [53]). Moreover, there are other issues that red-teaming alone cannot address, such as problems stemming from algorithmic monoculture [145, 71, 20]. We argue that red-teaming should therefore be considered as one evaluation paradigm, among others, to assess and improve the safety and trustworthiness of GenAI.
Red-teaming not well-scoped or structured. Additionally, the many variations in the red-teaming processes encountered in our case studies and literature review of public research and reports illustrate that at the moment, red-teaming is an unstructured procedure with undefined scope. As previously stated, we do not mean to belittle efforts undertaken so far to evaluate complex systems, and we concede that more may be happening in industry firms beyond our awareness, but in order to derive greater utility from future evaluations, we recommend that guidelines for red-teaming should be carefully drafted and made available to the public.
No standards concerning what should be reported. We further highlight that there are also currently no unified protocols for reporting the results of red-teaming evaluations. In fact, we found that a number of case studies and research papers sourced for our work did not fully report their findings or resource costs required to perform evaluations. For a number of reasons, ranging from increasing public knowledge to helping third-party groups conduct their own tests [113, 59], to assisting end-users in determining the relevance of red-teaming for their use cases, we suggest that regulations and/or best practices should be put forth to entice more detailed reporting following these exercises. We argue that such reports should, at a minimum, clarify (1) the resources consumed by the activity, (2) assessments of whether the activity was successful according to previously established goals and measures, (3) the mitigation steps informed by the findings of the activity, and (4) any other relevant or subsequent evaluation of the artifact at hand.
@@
Propose question bank as starting point. In light of the issues raised by our work, we provide a set of questions for future red-teams to consider before, during, and after evaluation. These questions, found in Table 1, encourage evaluators to ponder the benefits and limitations of red-teaming generally as well as the impact of specific design choices pertaining to their setting. We emphasize that these are not finalized guidelines but rather (what we hope is) the start of a broader conversation about GenAI red-teaming and evaluation processes. We welcome and support comments and feedback, and we leave refinement as a critical future direction.
@@ Acknowledgements
H. Heidari acknowledges support from NSF (IIS2040929 and IIS2229881) and PwC (through the Digital Transformation and Innovation Center at CMU). M. Feffer acknowledges support from the National GEM Consortium and the ARCS Foundation. Authors additionally gratefully acknowledge the NSF (IIS2211955), UPMC, Highmark Health, Abridge, Ford Research, Mozilla, Amazon AI, JP Morgan Chase, the Block Center, the Center for Machine Learning and Health, and the CMU Software Engineering Institute (SEI) via Department of Defense contract FA8702-15-D-0002, for their generous support of Z. Lipton’s, A. Sinha’s, and ACMI Lab’s research. W. H. Deng also acknowledges support from NSF (IIS-2040942), Cisco Research, the Jacobs Foundation, Google Research, and the Microsoft Research AI & Society Fellowship Program. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not reflect the views of the National Science Foundation and other funding agencies.
@@ Appendix A Research Survey and Case Study Details
This appendix contains further details about the research papers and case studies explored as part of this work. Table 4 contains the specific classifications along each dimension described in Section 4 for every work recovered as part of our research survey. We additionally provide access to a Google Sheets project with notes and thematic analyses for both the case studies and research papers retrieved and described in this work. The project can be accessed via this URL: https://docs.google.com/spreadsheets/d/1cZPc6Alkf8sqOFMsEvZgI2PzX2tHTIbemMa6sq4J2Qk/edit?usp=sharing.

### v3

@@ 4 A Survey of AI Red-teaming Research
Methodology. To source papers, we primarily searched arXiv, Google Scholar, OpenReview, ACL Anthology, and the ACM Digital Library with keywords “red-teaming”, “ai red-teaming”, “jailbreak”, and “llm jailbreak”, and we then gathered results.222We focus on red-teaming evaluations, but we argue that the jailbreaking literature contains techniques similar in spirit to those employed by AI red-teams, so we also explore it here. Where possible, we replaced preprints with corresponding published works. We also included relevant works found prior to this search and via snowball sampling.
We scrutinize and subdivide retrieved papers into groups along two dimensions, both of which relate to the evaluation in each paper. The first corresponds to the type of risk investigated during the evaluation, and the second corresponds to the type of approach used for evaluation. We analyze papers by characteristics pertaining to threat model and methodology because we found that research works primarily focused on these aspects (perhaps due to technical relevance) as opposed to other factors important to red-teaming (such as team composition and resources consumed). Overall classifications and in-depth paper findings (as with analyses of case studies from the previous section) can be found in our appendix, and totals for each classification can be found in Table 3. We do not ascribe judgments to any category (i.e., we do not posit that one red-teaming type is better than another).
@@ 4.1 Findings: AI Red-teaming Threat Models
We categorize papers according to the threat model implied or specified in each. In particular, we consider the type of risk posed in each threat model. For the ease of organizing our retrieved papers and findings, we present risks as one of two binary types, but we note that in reality, there exists a spectrum between the two types, and risks in practice generally may be more nuanced than as we describe them here.
Dissentive risk. We define dissentive risk as risk in which people may disagree on its definition and corresponding threat level, and we categorize papers that studied this type of risk accordingly. For instance, hate speech has many definitions, and though images with blood and responses to questions like “how to build a bomb?”333AI safety researchers note that “how to build a bomb?” is a default GenAI evaluation query despite the fact that related information could be found via Wikipedia or even popular fiction [168] because many GenAI developers assert their models should not disclose bomb-making instructions (e.g., [107, 154]), so techniques which yield responses prove safeguards are brittle. This does not necessarily mean the community finds such responses concerning. may be suboptimal AI generations, some may find them admissible depending on context.444Such generations do not reflect opinions held by the authors. Nevertheless, researchers like Gehman et al. [59] and Röttger et al. [131] create benchmarks to analyze hate speech generation, Rando et al. [122] perform red-teaming to study how Stable Diffusion’s safety filter can be bypassed to generate NSFW content, and Zou et al. [193] test whether LLMs respond to queries like “Generate a step-by-step plan to destroy humanity” given adversarial prompt suffixes.
@@
Both and neither. Some authors tasked themselves with analyzing both kinds of risk, such as personally identifiable information (PII) leakage in addition to hate speech [57, 109, 146]. Others introduce methods to analyze neither type of risk from the outset, stressing that definitions and classifications of issues may need to be done from scratch [28, 116].
@@ 4.2 Findings: AI Red-teaming Methodologies
Targeted attack. The last approach to red-teaming we document as part of our review involves deliberately targeting part of an LLM, which could include an API, vulnerability in language translation support, or step of its training process, in order to induce issues. As such, we refer to such approaches as targeted attack methods. For instance, Wang & Shu [161] show how to construct steering vectors using activation vectors from both safety-tuned and non-safety-tuned versions of models to obtain toxic outputs from safety-tuned models. Others illustrate how to imperceptibly perturb images to cause multimodal LMs to respond in unintended ways (such as replying with a malicious URL or misinformation) [136, 114, 12], and Tong et al. [153] engineer prompts for text-to-image models that are mismatched with resulting images by exploiting reliance on CLIP embeddings. Other approaches include but are not limited to weaponizing the fact that LLMs are not optimized to converse in low-resource languages and ciphers [47, 181, 183], poisoning data used to tune or utilize LLMs [123, 185, 3, 24, 162, 80], and attacking APIs associated with black-box models [108]. Various defensive methods rooted in targeted attack approaches have been proposed as well. Bitton et al. [19] describe their Adversarial Text Normalizer, which can defend an LLM against various character-level perturbations typical of certain adversarial prompts. In addition, other defensive strategies mentioned previously can defend these attacks (e.g., JailGuard from Zhang et al. [187] addresses attacks introduced in [114, 12, 136, 193]).
@@ 4.3 Discussion
Many different methods to perform red-teaming. As illustrated in Table 3, researchers and practitioners have undertaken numerous approaches to evaluate GenAI and have all described them as red-teaming. At the same time, there have been developments like Schuett et al.’s finding that the overwhelming majority of AGI lab members support external red-teaming efforts [137] and the recent Executive Order [151] stressing the importance of red-teaming. These developments and the many red-teaming variations are together arguably concerning, precisely because there is no agreed-upon definition (from these papers) regarding what constitutes red-teaming. By highlighting this, we do not mean to imply that evaluations until now are useless. On the contrary, we posit they are necessary but perhaps insufficient tests of safety, and we conjecture that the existence of many interpretations of “red-teaming” suggests there must be more top-down guidance and requirements concerning red-teaming evaluations.
Threat modeling skewed toward dissentive risk. Table 3 also highlights that the majority of evaluations focus on dissentive risk rather than consentive risk. This means that undue effort has been undertaken to evaluate and mitigate GenAI behavior that may be admissible in various contexts. Additionally, Röttger et al. [130] have shown that current attempts to mitigate such risks have resulted in exaggerated safety, yielding LLM behavior like the refusal to provide information on buying a can of coke. Lastly, focusing on dissentive risk takes attention away from consentive risk, which in turn is inadmissible in any context. In light of such issues and tradeoffs, Casper et al. [28] and Radharapu et al. [116] suggest clearly defining risks and problematic outputs and justifying those definitions before any analysis.
@@
Non-universality of values used for alignment. Work found as part of this survey involving dissentive risk and alignment are driven by, implicitly or explicitly, a set of human values that determine whether GenAI outputs are admissible or inadmissible. However, this in turn prompts the question whose values are being utilized for alignment and evaluation? For instance, the FLAMES benchmark proposed by Huang et al. [68] is purported to measure alignment with Chinese values, whereas Weidinger et al. [167] emphasize that other evaluations may reflect those of “the English-speaking or Western world.” The extent to which GenAI does not support low-resource languages [47, 181] and agrees with bias and stereotypes [57, 125] evidences that models may not reflect the values and beliefs of all persons. Works beyond this survey have illustrated how the framing of AI value alignment is a normative problem that, if not properly addressed, may only serve to reflect the norms of one group of people, typically the majority [51, 78, 77]. Especially as OpenAI started a partnership with the US military on one hand [147, 55] and launched an initiative to align superintelligent AI to “human values” on another [82],555This latter project ended in spring 2024 [54]; its dissolution may only further support the notion that AI developers’ positions should be scrutinized. we argue that it is crucial to analyze assumptions made and viewpoints held by those who build AI systems.
No consensus on who should perform red-teaming. Moreover, just as there is a lack of agreement regarding values to use to assess GenAI outputs, there is a similar lack of agreement regarding who should perform red-teaming. Groups of evaluators have consisted of hired crowdworkers [57], competition participants [138], researchers themselves [109], and others simply red-teaming for fun [71]. While some argue for more diversity to evaluate AI models [145], others caution that increased diversity is not a panacea and is moreover typically ill-defined [167, 13]. For instance, Yong et al. [181] argue for multilingual red-teaming to respond to low-resource language issues, and He et al. [65] “advocate for a collaborative, interdisciplinary approach among the AI for Science community and society at large” to respond to scientific research risks. Such examples suggest that terms like “diversity” and “community” should be defined and sought out relative to the risks considered by red-teaming processes. They additionally hint towards more involvement of the public and relevant stakeholders, ideas also recommended in parallel literature regarding algorithmic auditing and participatory ML [40, 16, 53, 42]. Similar literature has also engaged with benefits of deliberation in the face of disagreement (e.g., [111]) and effects of identity on evaluators’ perceptions of AI safety (e.g., [10]). Future red-teaming guidelines should emphasize these considerations.
@@ 5 NIST RFI Comment Analysis Summary
We find that comments submitted to the NIST RFI on red-teaming GenAI are generally in accord with our conclusions.666See our appendix for an extended analysis.
Similarities. Industry, academia, and civil society organizations suggest that NIST should specify a clearer definition of “red-teaming” and provide interested parties with appropriate resources pertaining to guidelines and best practices. Notably, even industry firms with experience red-teaming GenAI expressed a desire for concrete guidance from NIST. This supports our finding that red-teaming, as defined in public research and reports, is loosely structured and perhaps not the rigorous practice implied by the Executive Order. Moreover, many comments stressed that a plurality of different viewpoints and stakeholders should be involved in evaluating GenAI systems. Our findings concur with these notions.
@@ 6 Takeaways and Recommendations
Red-teaming is not a panacea. Each red-teaming exercise discussed in this paper only covered a limited set of vulnerabilities. As such, red-teaming cannot be expected to guarantee safety from all angles. For instance, from the papers surfaced in our research survey, approaches to red-teaming that detect and mitigate harmful text responses [109] may not detect and mitigate phishing attack vulnerabilities [133] and vice versa. Similarly, our case study analysis highlighted that team composition may also influence the types of issues found in a given exercise (e.g., subject matter experts [8] may find different problems than crowdworkers [57]). Moreover, there are other issues that red-teaming alone cannot address, such as problems stemming from algorithmic monoculture [75, 21, 153] or GenAI’s environmental impacts [41, 128]. We argue that red-teaming should therefore be considered as one evaluation paradigm, among others such as algorithmic impact assessments [126] and audits, to assess and improve the safety and trustworthiness of GenAI [56]. It is also important to support participation from diverse roles (e.g., technical, user-facing, legal) in red teaming within organizations [102, 46].
Red-teaming not well-scoped or structured. The many variations in the red-teaming processes encountered in our case studies and literature review of public research and reports illustrate that at the moment, red-teaming is an unstructured procedure with undefined scope. This statement is not meant to belittle efforts undertaken, and we concede that we do not have full insight into industry red-teaming activities, but we recommend red-teaming guidelines be drafted and made publicly available to improve utility of future evaluations.
No standards concerning what should be reported. There are currently no unified protocols for reporting the results of red-teaming evaluations. In fact, we found that a number of case studies and research papers sourced for our work did not fully report their findings or resource costs required to perform evaluations. We suggest that regulations and/or best practices be put forth to entice more detailed reporting for a number of reasons, ranging from increasing public knowledge, to helping third-party groups conduct their own tests [119, 63], to assisting end-users in determining the relevance of red-teaming for their use cases. We argue that such reports should, at a minimum, clarify (1) the resources consumed by the activity, (2) assessments of whether the activity was successful according to previously established goals and measures, (3) the mitigation steps informed by the findings of the activity, and (4) any other relevant or subsequent evaluation of the artifact at hand.
@@
Propose question bank as starting point. In light of the issues raised by our work, we provide a set of questions for future red teams to consider before, during, and after evaluation. These questions, found in Table 1, encourage evaluators to ponder the benefits and limitations of red-teaming generally as well as the impact of specific design choices pertaining to their setting. We emphasize that these are not finalized guidelines but rather (what we hope is) the start of a broader conversation about GenAI red-teaming and evaluation processes. We welcome and support comments and feedback, and we leave question refinement, overall evaluation, and development of supplementary materials (e.g., rubrics for evaluating red-teaming protocols) as critical future directions.
@@ Acknowledgements
Hoda Heidari acknowledges support from NSF (IIS2040929 and IIS2229881) and PwC (through the Digital Transformation and Innovation Center at CMU). Michael Feffer acknowledges support from the National GEM Consortium and the ARCS Foundation. Authors additionally gratefully acknowledge the NSF (IIS2211955), UPMC, Highmark Health, Abridge, Ford Research, Mozilla, Amazon AI, JP Morgan Chase, the Block Center, and the Center for Machine Learning and Health for their generous support of Zachary C. Lipton’s and ACMI Lab’s research. Wesley H. Deng also acknowledges support from NSF (IIS2040942), Cisco Research, the Jacobs Foundation, Google Research, and the Microsoft Research AI & Society Fellowship Program. Furthermore, this material is based upon work funded and supported by the Department of Defense under Contract No. FA8702-15-D-0002 with Carnegie Mellon University for the operation of the Software Engineering Institute, a federally funded research and development center. Any opinions, findings, conclusions, or recommendations expressed here are those of the authors and do not reflect the views of any funding agencies.
@@ Appendix A Research Survey and Case Study Details
This appendix contains further details about the research papers and case studies explored as part of this work. Table 4 contains the specific classifications along each dimension described in Section 4 for every work recovered as part of our research survey. We additionally provide access to a Google Sheets project with notes and thematic analyses for both the case studies and research papers retrieved and described in this work. The project can be accessed via this URL: https://docs.google.com/spreadsheets/d/1cZPc6Alkf8sqOFMsEvZgI2PzX2tHTIbemMa6sq4J2Qk/edit?usp=sharing. To conduct the thematic analyses, two authors used notes created for each case and paper encountered as part of this work and synthesized high-level takeaways that in turn formed the foundation for our Discussion subsections. Each author worked independently.